\begin{thebibliography}{14}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Bentivogli et~al.(2020)Bentivogli, Savoldi, Negri, Di~Gangi, Cattoni,
  and Turchi}]{bentivogli-etal-2020-gender}
Luisa Bentivogli, Beatrice Savoldi, Matteo Negri, Mattia~A. Di~Gangi, Roldano
  Cattoni, and Marco Turchi. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.619} {Gender in
  danger? evaluating speech translation technology on the {M}u{ST}-{SHE}
  corpus}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 6923--6933, Online. Association for
  Computational Linguistics.

\bibitem[{Cho et~al.(2019)Cho, Kim, Kim, and Kim}]{cho-etal-2019-measuring}
Won~Ik Cho, Ji~Won Kim, Seok~Min Kim, and Nam~Soo Kim. 2019.
\newblock \href {https://doi.org/10.18653/v1/W19-3824} {On measuring gender
  bias in translation of gender-neutral pronouns}.
\newblock In \emph{Proceedings of the First Workshop on Gender Bias in Natural
  Language Processing}, pages 173--181, Florence, Italy. Association for
  Computational Linguistics.

\bibitem[{Costa-juss{\`a} et~al.(2020)Costa-juss{\`a}, Li~Lin, and
  Espa{\~n}a-Bonet}]{costa-jussa-etal-2020-gebiotoolkit}
Marta~R. Costa-juss{\`a}, Pau Li~Lin, and Cristina Espa{\~n}a-Bonet. 2020.
\newblock \href {https://aclanthology.org/2020.lrec-1.502} {{G}e{B}io{T}oolkit:
  Automatic extraction of gender-balanced multilingual corpus of {W}ikipedia
  biographies}.
\newblock In \emph{Proceedings of the Twelfth Language Resources and Evaluation
  Conference}, pages 4081--4088, Marseille, France. European Language Resources
  Association.

\bibitem[{Currey et~al.(2022)Currey, Nadejde, Pappagari, Mayer, Lauly, Niu,
  Hsu, and Dinu}]{currey-etal-2022-mt}
Anna Currey, Maria Nadejde, Raghavendra~Reddy Pappagari, Mia Mayer, Stanislas
  Lauly, Xing Niu, Benjamin Hsu, and Georgiana Dinu. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.emnlp-main.288}
  {{MT}-{G}en{E}val: A counterfactual and contextual dataset for evaluating
  gender accuracy in machine translation}.
\newblock In \emph{Proceedings of the 2022 Conference on Empirical Methods in
  Natural Language Processing}, pages 4287--4299, Abu Dhabi, United Arab
  Emirates. Association for Computational Linguistics.

\bibitem[{Devinney et~al.(2022)Devinney, Bj\"{o}rklund, and
  Bj\"{o}rklund}]{10.1145/3531146.3534627}
Hannah Devinney, Jenny Bj\"{o}rklund, and Henrik Bj\"{o}rklund. 2022.
\newblock \href {https://doi.org/10.1145/3531146.3534627} {Theories of
  “gender” in nlp bias research}.
\newblock In \emph{Proceedings of the 2022 ACM Conference on Fairness,
  Accountability, and Transparency}, FAccT '22, page 2083–2102, New York, NY,
  USA. Association for Computing Machinery.

\bibitem[{Escud{\'e}~Font and
  Costa-juss{\`a}(2019)}]{escude-font-costa-jussa-2019-equalizing}
Joel Escud{\'e}~Font and Marta~R. Costa-juss{\`a}. 2019.
\newblock \href {https://doi.org/10.18653/v1/W19-3821} {Equalizing gender bias
  in neural machine translation with word embeddings techniques}.
\newblock In \emph{Proceedings of the First Workshop on Gender Bias in Natural
  Language Processing}, pages 147--154, Florence, Italy. Association for
  Computational Linguistics.

\bibitem[{Prates et~al.(2019)Prates, Avelar, and Lamb}]{prates2019assessing}
Marcelo O.~R. Prates, Pedro H.~C. Avelar, and Luis Lamb. 2019.
\newblock \href {http://arxiv.org/abs/1809.02208} {Assessing gender bias in
  machine translation -- a case study with google translate}.

\bibitem[{Ramesh et~al.(2022)Ramesh, Doddapaneni, Bheemaraj, Jobanputra, AK,
  Sharma, Sahoo, Diddee, J, Kakwani, Kumar, Pradeep, Nagaraj, Deepak, Raghavan,
  Kunchukuttan, Kumar, and Khapra}]{ramesh-etal-2022-samanantar}
Gowtham Ramesh, Sumanth Doddapaneni, Aravinth Bheemaraj, Mayank Jobanputra,
  Raghavan AK, Ajitesh Sharma, Sujit Sahoo, Harshita Diddee, Mahalakshmi J,
  Divyanshu Kakwani, Navneet Kumar, Aswin Pradeep, Srihari Nagaraj, Kumar
  Deepak, Vivek Raghavan, Anoop Kunchukuttan, Pratyush Kumar, and
  Mitesh~Shantadevi Khapra. 2022.
\newblock \href {https://doi.org/10.1162/tacl_a_00452} {Samanantar: The largest
  publicly available parallel corpora collection for 11 {I}ndic languages}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  10:145--162.

\bibitem[{Ramesh et~al.(2021)Ramesh, Gupta, and
  Singh}]{ramesh-etal-2021-evaluating}
Krithika Ramesh, Gauri Gupta, and Sanjay Singh. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.gebnlp-1.3} {Evaluating
  gender bias in {H}indi-{E}nglish machine translation}.
\newblock In \emph{Proceedings of the 3rd Workshop on Gender Bias in Natural
  Language Processing}, pages 16--23, Online. Association for Computational
  Linguistics.

\bibitem[{Rarrick et~al.(2023)Rarrick, Naik, Mathur, Poudel, and
  Chowdhary}]{rarrick2023gate}
Spencer Rarrick, Ranjita Naik, Varun Mathur, Sundar Poudel, and Vishal
  Chowdhary. 2023.
\newblock \href {http://arxiv.org/abs/2303.03975} {Gate: A challenge set for
  gender-ambiguous translation examples}.

\bibitem[{Saunders et~al.(2020)Saunders, Sallis, and
  Byrne}]{saunders-etal-2020-neural}
Danielle Saunders, Rosie Sallis, and Bill Byrne. 2020.
\newblock \href {https://aclanthology.org/2020.gebnlp-1.4} {Neural machine
  translation doesn{'}t translate gender coreference right unless you make it}.
\newblock In \emph{Proceedings of the Second Workshop on Gender Bias in Natural
  Language Processing}, pages 35--43, Barcelona, Spain (Online). Association
  for Computational Linguistics.

\bibitem[{Savoldi et~al.(2021)Savoldi, Gaido, Bentivogli, Negri, and
  Turchi}]{savoldi-etal-2021-gender}
Beatrice Savoldi, Marco Gaido, Luisa Bentivogli, Matteo Negri, and Marco
  Turchi. 2021.
\newblock \href {https://doi.org/10.1162/tacl_a_00401} {Gender bias in machine
  translation}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  9:845--874.

\bibitem[{Stanovsky et~al.(2019)Stanovsky, Smith, and
  Zettlemoyer}]{stanovsky-etal-2019-evaluating}
Gabriel Stanovsky, Noah~A. Smith, and Luke Zettlemoyer. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1164} {Evaluating gender bias
  in machine translation}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 1679--1684, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Zhao et~al.(2018)Zhao, Wang, Yatskar, Ordonez, and
  Chang}]{zhao-etal-2018-gender}
Jieyu Zhao, Tianlu Wang, Mark Yatskar, Vicente Ordonez, and Kai-Wei Chang.
  2018.
\newblock \href {https://doi.org/10.18653/v1/N18-2003} {Gender bias in
  coreference resolution: Evaluation and debiasing methods}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 2 (Short Papers)}, pages 15--20, New Orleans, Louisiana.
  Association for Computational Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[]{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage{devanagari}
\usepackage{graphicx}
\usepackage[table]{colortbl}
\usepackage{booktabs}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{Don't Overlook the Grammatical Gender: Bias Evaluation for Hindi-English Machine Translation}

\author{Pushpdeep Singh \\
  National Institute of Technology, Hamirpur \\ Anu, Hamirpur, India \\
  \texttt{pushpdeep30@gmail.com} \\}

\begin{document}
\maketitle
\begin{abstract}
Neural Machine Translation (NMT) models, though state-of-the-art for translation, often reflect social biases, particularly gender bias. Existing evaluation benchmarks primarily focus on English as the source language of translation. For source languages other than English, studies often employ gender-neutral sentences for bias evaluation, whereas real-world sentences frequently contain gender information in different forms. Therefore, it makes more sense to evaluate for bias using such source sentences to determine if NMT models can discern gender from the grammatical gender cues rather than relying on biased associations. To illustrate this, we create two gender-specific sentence sets in Hindi to automatically evaluate gender bias in various Hindi-English (HI-EN) NMT systems. We emphasise the significance of tailoring bias evaluation testsets to account for grammatical gender markers in the source language. 

\end{abstract}

\section{Introduction}

State-of-the-art NMT models develop gender bias \citep{prates2019assessing} during their training by picking up spurious correlations in their training data. Due to their poor coreference resolution ability, they often rely on biased associations with, say, occupation terms to disambiguate the gender of pronouns (\citealp{stanovsky-etal-2019-evaluating, saunders-etal-2020-neural}). Also, in many cases, these models translate gender-neutral sentences into gendered ones (\citealp{prates2019assessing, cho-etal-2019-measuring}) especially giving a `masculine default' translation or one which is based on a stereotype. 
This problem also exists for HI-EN Machine Translation \citep{ramesh-etal-2021-evaluating}. When put to use, such systems can cause various harms \citep{savoldi-etal-2021-gender}. Thus, evaluating and mitigating such biases from NMT models is critical to ensure fairness. 

Prior research evaluating gender bias in machine translation has predominantly centered around English as the source language (\citealp{stanovsky-etal-2019-evaluating, bentivogli-etal-2020-gender, costa-jussa-etal-2020-gebiotoolkit, currey-etal-2022-mt, rarrick2023gate}). However, these evaluation methods or testsets don't seamlessly extend to other source languages, especially the ones with grammatical gender. For instance, in Hindi, elements like pronouns, adjectives, and verbs are inflected with gender. Nonetheless, prior studies in other source languages often utilise gender-neutral sentences for bias evaluation using TGBI score (\citealp{cho-etal-2019-measuring, ramesh-etal-2021-evaluating}). In our experiments with TGBI Evaluation, we found that it doesn't expose gender bias present in various NMT models.

Therefore, in this work, we propose to evaluate NMT models for bias using sentences with grammatical gender cues of the source language. This allows us to ascertain whether NMT models can discern the accurate gender from context or if they depend on biased correlations. Using Hindi as the source language in NMT, we develop two bias evaluation benchmarks: \emph{OTSC-Hindi} and \emph{WinoMT-Hindi} and evaluate various HI-EN NMT models for gender bias. We highlight the importance of creating such benchmarks for source languages with expressive gender markers.

\begin{table*}[tbh!]
  \centering
  \begin{tabular}{@{}lcccccccc@{}}
    \toprule
    & \multicolumn{2}{c}{\textbf{IT}}                                                                & \multicolumn{2}{c}{\textbf{GT}}                                                                & \multicolumn{2}{c}{\textbf{MS}}                                                                 & \multicolumn{2}{c}{\textbf{AWS}}                                                             \\
\multicolumn{1}{c}{\textbf{Sentence Set}}    & \multicolumn{1}{c}{$p_m$} & \multicolumn{1}{c}{$p_f$} & \multicolumn{1}{c}{$p_m$} & \multicolumn{1}{c}{$p_f$} & \multicolumn{1}{c}{$p_m$} & \multicolumn{1}{c}{$p_f$}  & \multicolumn{1}{c}{$p_m$} & \multicolumn{1}{c}{$p_f$} \\ \midrule
    \multicolumn{1}{l|}{\textit{\small{Female Speaker, Female Friend}}} & \textbf{98.41}                                                  & \multicolumn{1}{c|}{1.59$^*$}      & 1.68                         & \multicolumn{1}{c|}{\textbf{98.32$^*$}}      & \textbf{98.97}                                   & \multicolumn{1}{c|}{1.03$^*$}      & \textbf{95.61}                                                    & 4.39$^*$                           \\
    \multicolumn{1}{l|}{\textit{\small{Female Speaker, Male Friend}}} & {\textbf{99.25$^*$}}                                  & \multicolumn{1}{c|}{0.75}      & \textbf{90.66$^*$}  & \multicolumn{1}{c|}{9.34}      & \textbf{99.72$^*$}                                                     & \multicolumn{1}{c|}{0.28}      & \textbf{95.70$^*$}                                              & 4.30                            \\
    \multicolumn{1}{l|}{\textit{\small{Male Speaker, Female Friend}}} & \textbf{99.35}                                                   & \multicolumn{1}{c|}{0.65$^*$}      & 2.43                                & \multicolumn{1}{c|}{\textbf{97.57$^*$}}      & \textbf{66.01}                                       & \multicolumn{1}{c|}{33.99$^*$}      & \textbf{99.29}                                                    & 2.71$^*$                            \\
        \multicolumn{1}{l|}{\textit{\small{Male Speaker, Male Friend}}} & \textbf{99.91$^*$}                                                      & \multicolumn{1}{c|}{0.09}      & \textbf{96.45$^*$}                               & \multicolumn{1}{c|}{3.55}      & \textbf{98.60$^*$}                                          & \multicolumn{1}{c|}{1.40}      & \textbf{97.48$^*$}                                               & 2.52                            \\\bottomrule
    \end{tabular}%

  \caption{
    Evaluation using the OTSC-Hindi test set. Here, $p_m$ and $p_f$ are the percentages of sentences translated as male and female for the speaker’s friend. $*$ corresponds to the percentage of sentences translated into the true label for each sentence set. Bold values indicate the maximum percentage of sentences translated into a single gender class.}
    \label{tab:results_otsc}
\end{table*}
\begin{table}[tbh!]
  \centering
  \begin{tabular}{@{}lcccc@{}}
    \toprule
    & \multicolumn{1}{c}{\textbf{$Acc$}} & \multicolumn{1}{c}{\textbf{$\Delta_G$}} & \multicolumn{1}{c}{\textbf{$\Delta_S$}} & \multicolumn{1}{c}{\textbf{$N$}} \\ \midrule
    \textit{\small{\textbf{IndicTrans}}} &48.9 &48.5 &-0.1$^{\bullet}$ &6.2    \\
    \textit{\small{\textbf{Google Translate}}} &69.0$^{\star}$ &10.6$^{\diamond}$ &-3.8 &5.3     \\ 
    \textit{\small{\textbf{Microsoft Translator}}} &57.7 &32.9 &0.2$^{\bullet}$ &4.1     \\ 
    \textit{\small{\textbf{AWS Translate}}} &49.9 &51.9 &-0.2$^{\bullet}$ &2.8     \\ 
    \bottomrule
    \end{tabular}%
  \caption{Comparison of performance of various NMT Models on WinoMT-Hindi. ${\star}$ indicates significantly highest value, ${\diamond}$ indicates significantly lowest value, ${\bullet}$ indicates near about values for \emph Acc, $\Delta_G$ and $\Delta_S$, respectively. \emph{N} represents percentage of sentences translated into gender-neutral form. }
  \label{tab:results-winomt}
  \end{table}

\section{Experimental Setup}

We create two bias evaluation testsets in Hindi to evaluate gender bias: Occupation Testset with Simple Context (\textit{OTSC-Hindi}) and \textit{WinoMT-Hindi}, which account for diverse grammatical gender cues in Hindi. Using these, we evaluate four HI-EN NMT Models: IndicTrans (IT) \citep{ramesh-etal-2022-samanantar}, Google Translate (GT), Microsoft Translator (MS) and AWS Translate (AWS). Previous work by \citet{ramesh-etal-2021-evaluating} focused only on the gender-neutral side of Hindi; however, we demonstrate how we can account for gender-specified context for better evaluation of bias. Code and data are publicly available\footnote{\url{https://github.com/iampushpdeep/Gender-Bias-Hi-En-Eval}}.

\section{OTSC-Hindi}

Similar to \citet{escude-font-costa-jussa-2019-equalizing}, we create a Hindi version with grammatical gender cues: `` {\dn m\4{\qva} us\? kAPF smy s\?} \{{\dn jAntA{\rs ,\re}jAntF}\} {\dn \8{h}\1{\rs,\re}}\{{\dn m\?rA{\rs,\re}m\?rF}\} {\dn do-t} \textbf{[occupation]} {\dn kA kAm} \{{\dn krtA{\rs ,\re}krtF}\} {\dn h\4.} ''(\textit{I have known [him/her] for a long time, my friend works as a [occupation].}) Unlike the English version, this template specifies the gender of the speaker using a gender-inflected verb, i.e. ``{\dn jAntA}(\textit{jaanta})''(m) and ``{\dn jAntF}(\textit{jaanti})''(f). The possessive pronoun is also gender inflected based on the gender of the speaker's friend: ``{\dn m\?rA}(\textit{mera})''(m) and ``{\dn m\?rF}(\textit{meri})''(f) which decides the use of verb ``{\dn krtA}(\textit{karta})''(m) and ``{\dn krtF}(\textit{karti})'' (f).  However, the English translation only specifies the gender of speaker's friend. Based on four possibilities in the template, we construct these four sets of sentences.

\section{WinoMT-Hindi}
\citet{stanovsky-etal-2019-evaluating} composed a challenge set called \textit{WinoMT} for evaluating gender bias in NMT models. We contextualize this test set for the evaluation of bias in HI-EN Translation by manually creating ``WinoMT-Hindi'', which consists of 704 WinoBias-like sentences \cite{zhao-etal-2018-gender} in Hindi but modified to include gender cues of the language, mainly: gender-inflected adjectives, postpositions, and verbs. We can mark the gender of the target (English) by simply checking for the presence of male pronouns (he, him or his) or female pronouns (she or her) in the translation. Interestingly, we also observe that a few sentences are translated into a gender-neutral form. For gender bias evaluation, we use the metrics \emph{Acc}, $\Delta_G$ and $\Delta_S$ given by \citet{stanovsky-etal-2019-evaluating}.

\section{Evaluation Results}
The results are shown in Table~\ref{tab:results_otsc} and Table~\ref{tab:results-winomt}. Based on these results, most models show heavy bias against the female gender and prefer masculine default in translations. On WinoMT-Hindi, these models perform only as good as random guessing, while Google Translate performs better at discerning gender using the grammatical gender cues in the source. On the other hand, evaluation based on TGBI (results not shown here) gives almost similar performance on all NMT systems which shows that it is not good at exposing gender bias.

\section{Conclusion and Future Work}
We highlighted the need for contextualising NMT bias evaluation for non-English source languages. In future, we plan to extend our evaluation beyond following a particular template and be accommodating to all gender identities \citep{10.1145/3531146.3534627}. 

\bibliography{anthology,custom}

\end{document}